# Assessing Physics Students’ Scientific Argumentation using Natural Language Processing

### v1

@@ 4.1 Dataset
This study was implemented in a first-semester calculus-based physics class for future engineers, at a large U.S. Midwestern land grant University. The annual enrollment at the time of the study is approximately 2500 students (1100 in fall and 1400 in spring). The course is built around three key principles: momentum, energy, and angular momentum, and it follows Chabay and Sherwood’s Matter & Interactions [47]. The course has three components: lectures (two 50-minute-long session), lab (one 110-minute-long session, and recitations (one 50-minute-long session). The recitations are the context of this study.
The recitations are led by one graduate teaching assistant (GTA) with the aid of an undergraduate teaching assistants (UTA). During each Recitation session, the GTA’s spend a short time at the beginning of the recitation introducing the problem and relevant information. They start off each section by going over a provided Powerpoint presentation introducing the recitation problem. Each presentation takes around 5 to 10 minutes. In earlier weeks of the semesters, students receive more help on starting the problem, such as going over system and surroundings as a class. As the semesters progress students receive more general help on understanding the problem and less help on starting the practical aspects of it. Then students are expected to work together in groups of 3-5 at their table to solve the problem. They often use a whiteboard to create a collaborative workspace to discuss and share their work. As students work together, the GTA and UTA are available and walking around the room to help groups when they have questions. Even though they work in groups, every student is expected to provide their answers to questions in an individual Jupyter Notebook file (.ipynb extension). They edit this file on Google Colab in class. Students discuss the questions collaboratively and enter their answers to these questions and input an image of their written work.
After they are finished, they are expected to submit the file and a PDF of the file into Brightspace. The notebook file is uploaded so we have access to easily extract student responses. A PDF is uploaded for grading purposes. Students are graded as a group by GTAs on a rubric provided. GTAs choose one student randomly from each group, grade them for correctness, and assign everyone in the group this grade. Students are not expected to spend much, if any, time on the recitation outside of the 50-minute section, so there should be very little individual work on the recitation other than expressing their answers in their own words. Due to the nature of the grading, students are encouraged to ask the TAs questions and make sure everyone in the group is on the same page before they leave the Recitations session.
@@
4.2 Scaffolding
Throughout the last two years, students were tasked with writing argumentation essays at the end of most recitations. Table 1 outlines how the scaffolds progressed each semester. The example question in the second row of each column represents the question students were asked at the end of the semester after receiving their full set of scaffolds. The dataset we report on is in response to question in the table.
The study started in Fall 2022 when students received no instruction on argumentation, except of the prompt shown in Table 1. In Spring 2023, we implemented argumentation prompts based upon McNeil and Krajcik’s model within the recitation. We gave students the definitions of claim, evidence, and reasoning (CER) in pieces as the semester progressed, while prompting them to list their own CER in parts before writing a full argument of their own. Finally in Fall 2023 and Spring 2024 students were given similar scaffolds to that of Spring 2023. However, instead of immediately having students list their own CER, we gave students statements and had students answer which statements were CER. Then in the middle portion of the semester students listed their own CER. Finally they constructed their own arguments at the end of the semester.
Table 1: Argumentation Scaffolding
@@ 4.2 Scaffolding
Module 1 Students were asked to construct arguments Students were given definitions of Claim, Evidence, and Reasoning (CER) and asked to list their CER, separately. Students were given definitions of CER and prompts to identify CER statements.
Module 2 Students were asked to construct arguments Students were given definitions of CER and asked to construct arguments. Students were given definitions of CER and asked to list their CER, separately.
Module 3 Students were asked to construct arguments Students were given definitions of CER and asked to construct arguments. Students were given definitions of CER and asked to construct arguments.
@@ 4.3 Topic Modeling
We analyzed open-ended student scientific arguments from each of the four semesters separately. Although, students were given instruction to use complete only words and complete sentences to construct their essays, there still required cleaning of the text. We began by employing a comprehensive text pre-processing pipeline to clean student argumentation essays. This included punctuation and number removal, conversion to lowercase, and stopword (e.g. ’the’, ’and’ , ’as’, etc.) filtering using NLTK [48]. Additionally, we implemented automated spell checking using the pyspellchecker library to correct common typos. Responses with fewer than 10 words were filtered out to reduce noise and trivial responses.
We applied non-negative matrix factorization (NMF) to extract latent topics from the student essays. First, the pre-processed text was vectorized using the TF-IDF method, capturing term importance across the corpus. We then applied the NMF algorithm [46] with ten components (topics). Each component represents a distribution over words, and each essay is represented as a mixture of topics.
To identify the optimal number of topics for NMF, we evaluated both the reconstruction error [46] and coherence scores across topic counts ranging from 2 to 30. The reconstruction error, which measures the discrepancy between the original TF-IDF matrix and its reconstruction from the NMF model, was calculated for each topic configuration. This metric is used to assess how well the NMF model approximates the original data. Coherence scores provide an indication of how interpretable the topics are. We used the Coherence Model [49] from the gensim library to compute coherence scores for the topics generated by the NMF model. As expected, reconstruction error decreased monotonically with additional topics, Figure 2, reflecting better matrix approximation. However, coherence scores, which measure the semantic interpretability of topics, fluctuated and did not exhibit a clear maximum . Based on a balance between interpretability and parsimony, and in line with prior work using similar data, we selected 10 topics as a reasonable compromise for downstream analysis.
@@
Figure 2: NMF Reconstruction Error vs Number of Topics
To interpret the topics, we examined the top-weighted words associated with each component. Each essay was assigned to its most dominant topic using the argmax of the topic mixture vector. We visualized topic distribution across the corpus using a histogram, highlighting the number of student essays per topic. To compare the semesters, we analyzed the representative words of each topic, the number of essays in each topic, and performed a concentration analysis [50] of each semester. Bao and Redish [50] introduced this method to measure how students’ responses on multiple-choice questions are distributed. We have applied this method to measure how our student essays are distributed over different topics.
5 Results and Discussion
We report on the top five words of each topic and the essay distribution over topics of each semester. The resulting top five words were calculated to assess what each topic represented (2,3,4,5), and the number of essays in each topic are represented in the histograms (3(a),3(b),3(c), 3(d)) shown below.
Topics 1 2 3 4 5 Counts
Topic 0 energy kinetic change work equal 96
Topic 1 acceleration torque net angular force 73
Topic 2 system extended point particle part 80
Topic 3 solution correct made us sure 56
Topic 4 force applied cylinder work distance 90
Topic 5 mass center velocity force solve 64
Topic 6 used solved part energy principle 60
Topic 7 air resistance problem assumed cylinder 66
Topic 8 equation using part solve theorem 105
Topic 9 find velocity use final principle 51
Table 2: Top 5 words for each topic in Fall 2022 using NMF
Topics 1 2 3 4 5 Counts
Topic 0 energy kinetic translational rotational potential 71
Topic 1 cylinder incline slipping without rolling 84
Topic 2 used part principle find momentum 88
Topic 3 force friction applied distance gravity 98
Topic 4 system point particle extended wheel 66
Topic 5 mass center velocity find using 59
Topic 6 work change done energy equal 58
Topic 7 equation acceleration answer get vim 86
Topic 8 problem evidence claim earth surroundings 90
Topic 9 use principle solve energy need 65
Table 3: Top 5 words for each topic in Spring 2023 using NMF
Topics 1 2 3 4 5 Counts
Topic 0 system extended particle point part 68
Topic 1 energy kinetic work done equal 170
Topic 2 acceleration momentum angular linear torque 80
Topic 3 force applied distance twice friction 83
Topic 4 use energy principle problem must 87
Topic 5 friction work contact point move 72
Topic 6 object pulling problem would large 45
Topic 7 disk upwards slipping force counteracts 18
Topic 8 cylinder ramp slipping belt earth 96
Topic 9 used find velocity equation mass 113
Table 4: Top 5 words for each topic in Fall 2023 using NMF
Topics 1 2 3 4 5 Counts
Topic 0 system particle point extended friction 87
Topic 1 cylinder slipping without incline force 137
Topic 2 energy kinetic work change done 178
Topic 3 momentum angular principle find use 73
Topic 4 ramp surroundings earth belt system 94
Topic 5 force applied friction equal times 116
Topic 6 velocity mass center find using 78
Topic 7 work displacement ball slip slipping 47
Topic 8 cylinders speed workenergy theorem motion 39
Topic 9 solve part principle energy used 124
Table 5: Top 5 words for each topic in Spring 2024 using NMF
In fall 2022, we found the top five words for each topic and the number of essays allocated to each topic. Within the topics there were some repetition of words between topics: "force" and "part" were repeated three times while "energy", "work", "cylinder", "velocity", "solve", and "principle" were repeated twice. Across the topics, there were three relative peaks in Topics 0, 4, and 8 with the largest peak at Topic 8. Topic 0 focuses on concepts from the energy principle such as "energy", "kinetic, and "work. Topic 4 focuses on superficial features such as cylinder", "applied", "distance" and ideas of work such as "work" and "force". Topic 8 focuses on apparent mathematical steps to solve the problem which is apparent with the words "equation", "using", "part", "solve", and "theorem". Even though there are three peaks, the topics seem to be evenly distributed, meaning there is a broad range of topics that students focused on equally.
In Spring 2023, we found the top five words for each topic and the number of essays allocated to each topic. Within the topics there were some repetition of words between topics: "energy" was repeated 3 times while "find" and "principle" were repeated twice. Across the topics, there were two relative peaks in Topics 3 and 8 with the largest peak at Topic 3. Topic 3 focuses primarily on the forces such as "friction", "gravity", and literally "force". "Applied" likely is in reference to applied force. Topic 8 focuses not on physics concepts but argumentation terms such as "evidence", "claim, and generic "problem". Even though there are two peaks, the topics seem to be evenly distributed, meaning there is a broad range of topics students equally focused on.
In fall 2023, we found the top five words for each topic and the number of essays allocated to each topic. Within the topics there were some repetition of words between topics: "energy", "point", "work, "force", "friction", "problem" and "slipping" were repeated twice. Across the topics, there was one relative peak Topic 1. Across the topics, there was one relative peak in Topic 1, which included 170 essays. Topic 9 had the second most number of essays at 113. Topic 1 focuses on words we would expect from the energy principle, such as "energy", "kinetic", and "work". In this semester, there seems to be an unequal distribution of topics with essays skewing towards Topic 1.
In spring 2024, we found the top five words for each topic and the number of essays allocated to each topic. Within the topics there were some repetition of words between topics: "energy", "system, "friction", "slipping", "work", "force", "principle", and "find" were repeated twice. One odd outlier that appeared was a conjunction of two words "workenergy". Across the topics, there was one relative peak Topic 2, which included 178 essays. Topic 1 had the second most number of essays at 137. Much like in fall 2023, topic 2 focuses on words we would expect from the energy principle, such as "energy", "kinetic", and "work". In this semester, there seems to be an unequal distribution of topics with essays skewing towards Topic 2.
(a) Fall 2022
(b) Spring 2023
(c) Fall 2023
(d) Spring 2024
Figure 3: Number of Essays in Each NMF Topic across Four Semesters
After investigating the top words and distribution, we found the first two semesters, which had either no or less scaffolding, seemed to be evenly distributed over topics. Which means as a whole we had students focusing on important and correct concepts, such as energy, but equally focusing on trivially concepts, such as mathematical steps. In the last two semesters, in which students received the most scaffolding, we saw unequal distributions, with peaks in topics that focused on energy ideas. This indicates that students in these semesters tended to focus more on physics oriented ideas.
@@ 5 Results and Discussion
To further investigate the distribution across topics, we performed a concentration analysis and report on the coherence factor in each semester in Figure 4. We see an increase in concentration from Fall 2022 to Spring 2024. The Fall 2023 and Spring 2024 semesters have higher concentrations of essays than Fall 2022 and Spring 2023 semesters. It is interesting to note we may begin to see a pattern emerging between Fall and Spring semesters. The concentration of Spring 2023 is less than that of Fall 2022, and the concentration of Spring 2024 is less than that of Fall 2023. More investigation needs to be done to determine if this pattern holds for more semesters; however, there could be interesting implications to this finding. When students have misconceptions it tends to result in more diverse incorrect answers, while correct student answer tend to reflect more uniform understanding [51]. Therefore, in Spring semesters students may have more misconceptions than students in fall semesters, since we see less coherence.

### v2

@@ 3.1 Dataset
This study was implemented in a first-semester calculus-based physics course for future engineers, at a large U.S. Midwestern land grant University. The annual enrollment in the course at the time of the study is approximately 2500 students (1100 in fall and 1400 in spring). The course is built around three key principles: momentum, energy, and angular momentum, and it follows Chabay and Sherwood’s Matter & Interactions [chabay2011matter]. The course has three components: lectures (two 50-minute-long session), lab (one 110-minute-long session, and recitations (one 50-minute-long session). The recitations are the context of this study.
The recitations are led by one graduate teaching assistant (GTA) with the aid of an undergraduate teaching assistants (UTA). During each Recitation session, the GTA’s spend a short time at the beginning of the recitation introducing the problem and relevant information. They start off each section by going over a provided Powerpoint presentation introducing the recitation problem. Each presentation takes around 5 to 10 minutes. In earlier weeks of the semesters, students receive more help on both the conceptual aspects of the problem (i.e. identifying the relevant principles and concept) as well as on the procedural aspects (i.e. selecting the appropriate system and surroundings). Students also receive scaffolding on the argumentation aspects of the problem, which is described below in more detail. As the semester progresses students continue to receive help on the conceptual aspects of the problem, but less help on starting the procedural aspects of it. Then students are expected to work together in groups of 3-5 at their table to solve the problem. They often use a whiteboard to create a collaborative workspace to discuss and share their work. As students work together, the GTA and UTA are available and walking around the room to help groups when they have questions. Even though they work in groups, every student is expected to provide their answers to questions in an individual Jupyter Notebook file (.ipynb extension). They edit this file on Google Colab in class. Students discuss the questions collaboratively and enter their answers to these questions and input an image of their written work.
After they are finished, they are expected to submit the file and a PDF of the file into the Learning Management System (Brightspace). The notebook file is uploaded so we have access to easily extract student responses. A PDF is uploaded for grading purposes. Students are graded as a group by GTAs on a rubric provided. GTAs choose one student randomly from each group, grade their uploaded recitation file based on a rubric, and assign everyone in the group this grade. Students are not expected to spend much, if any, time on the recitation outside of the 50-minute section, so there should be very little individual work on the recitation other than expressing their answers in their own words. Due to the nature of the grading, students are encouraged to ask the TAs questions and make sure everyone in the group is on the same page before they leave the recitation session.
@@
3.2 Scaffolding Scientific Argumentation
Table 1: Summary of recitation scaffolds by semester.
@@ 3.2 Scaffolding Scientific Argumentation
Module 1 Students were asked to construct arguments. Students were given definitions of CER and asked to list their CER separately. Students were given definitions of CER and prompts to identify CER statements.
Module 2 Students were asked to construct arguments. Students were given definitions of CER and asked to construct arguments. Students were given definitions of CER and asked to list their CER separately.
Module 3 Students were asked to construct arguments. Students were given definitions of CER and asked to construct arguments. Students were given definitions of CER and asked to construct arguments.
@@
Throughout the last four semesters (Fall 2022 through Spring 2024), students were tasked with writing argumentation essays at the end of most recitations. Table 1 outlines how the scaffolds progressed each semester. The example question in the second row of each column represents the question students were asked at the end of the semester after receiving their full set of scaffolds. The dataset we report on is in response to question in Fig. 1.
The study started in Fall 2022 when students received no instruction on how to construct an argument; they were given the prompt shown in Table 1. In Spring 2023, we implemented argumentation prompts based upon McNeil and Krajcik’s model [mcneill2010scientific] within the recitation. We gave students the definitions of claim, evidence, and reasoning (CER):
  • A CLAIM is every decision you took to solve the problem i.e. choice of principle/concept, system/surrounding, approximation/assumption, or key steps you used in your problem-solving strategy – It is an assertion that that the decision was the right one to answer the key question(s) in the problem statement.
  • EVIDENCE is the information either provided in the problem or something that you know (e.g. from the physics class) to be factually correct that helped you make the decision above.
  • REASONING is an explanation of why the EVIDENCE supports the CLAIM.
We provided the definitions starting with claims as the semester progressed, while prompting students to list their own CER in parts before writing a full argument of their own. Finally in Fall 2023 and Spring 2024 students were given similar scaffolds to that of Spring 2023. However, instead of immediately having students list their own CER, we gave students statements and had students label which statements were claim, evidence, or reasoning. Then in the middle portion of the semester students listed their own CER statements. Finally they constructed their own arguments at the end of the semester. The breakdown of these scaffolds and the final argumentation question asked to students are listed in Table 1.
@@ 3.3 Topic Modeling
We analyzed open-ended student scientific arguments from each of the four semesters separately. Although, students were given instructions to use only words and complete sentences to construct their essays, the data still required significant cleaning of the text. We thus began by employing a comprehensive text pre-processing pipeline to clean student argumentation essays. This included punctuation and number removal, conversion to lowercase, and removal of stopwords (e.g. ’the’, ’and’ , ’as’, etc.) filtering using the Natural Language Toolkit (NLTK) [bird2009natural]. Additionally, we implemented automated spell checking using the PySpellChecker [pyspellchecker] library to correct common typos. Responses with fewer than ten words were filtered out to reduce noise.
We applied Non-negative Matrix Factorization (NMF) to extract latent topics from the student essays. First, the pre-processed text was vectorized using the TF-IDF method, capturing term importance across the corpus. We then applied the NMF algorithm [lee1999learning] with ten components (topics). Each component represents a distribution over words, and each essay is represented as a mixture of topics.
One common method to determine the optimal number of topics is known as the "elbow method" where you plot a model fit, such as reconstruction error[lee1999learning] or coherence scores [coherence2015], across a number of topics. To identify the optimal number of topics for NMF, we plotted both the reconstruction error [lee1999learning] and coherence scores [coherence2015] across topic counts ranging from two to thirty. The reconstruction error, which measures the discrepancy between the original TF-IDF matrix and its reconstruction from the NMF model, was calculated for each topic configuration. This metric is used to assess how well the NMF model approximates the original data. Though the reconstruction error decreased monotonically with additional topics reflecting better matrix approximation a definitive "elbow" did not emerge for any semester. For the coherence scores, which measure the semantic interpretability of topics, fluctuated and did not exhibit a clear maximum. Since there was no clear quantitative metric to identify ideal number of topics, we selected 10 topics in line with prior work [bralin2023analysis] using similar data context.
3.4 Interpretation of Topics
To compare the semesters, we analyzed the representative words of each topic, the number of essays in each topic, and the overall distribution of student responses across topics using a concentration analysis [bao2001concentration] of each semester. To interpret the topics across semesters, we examined the top-weighted words associated with each component from the NMF model. Each essay was assigned to the topic with the highest topic weight. We then visualized topic distribution across the corpus using histograms that highlight the number of student essays per topic.
The concentration analysis introduced by Bao and Redish [bao2001concentration] was originally developed to measure how students’ responses on multiple-choice questions are distributed. Since we are interested in measuring student membership to a topic, this analysis will apply a novel insight into our student distribution. In this study, we have applied this method to quantify how student essays are distributed over different topics. The concentration factor, $C$, is calculated as:
4 Results
@@
Figure 2: Representative words and distribution of essays in NMF topics across scaffolded semesters
Our goal is to utilize topic modeling to determine if we can capture the longitudinal change in the broad themes of student arguments throughout the semester. To achieve this we report on the top five words of each topic, exemplar essays from relevant topics, and the essay distribution over topics of each semester.
The resulting top five words were calculated to assess what each topic represented, shown in Figure 2. The number of essays in each topic are represented in the bar graphs of Figures 2(a), 2(b), 2(c), and 2(d) shown below. We strove to improve students argumentation with the scaffolds in such a way that students focused more on the underlying physics principles of the problem as opposed to the surface features or mathematical steps of solving the problem. For this specific problem, a correct solution would focus on using the Energy Principle: $\Delta E_{system}=W_{surroundings}$ and concepts involving energy. Therefore, we would ideally see the top words of the topics focused on types of energy and/or work.
First, we looked at Fall 2022 topic distribution and top words in Figure 2(a). We found the topics to be fairly evenly distributed with minor peaks at Topics 0, 4, and 8. Our overall takeaway from these representative words is that there was some focus on explicit physics principles (i.e. Topics 0 and 4), which became apparent from emergent words such as "energy", "kinetic", and "force". There was much more focus on mathematical steps (i.e. Topics 3, 6, 7, 8, and 9) shown from the representative words, such as "solution", "used", "equation" and "find". To further look into the topics we report on a representative essay from the topic with the highest student membership. For Fall 2022, this would be Topic 8. Each student essay is comprised of variations of topics. For an essay to be categorized in a certain topic the essay must have a majority percentage of that topic. In Topic 8, for example, one of the most representative essays is classified as 84.41 % of Topic 8, 11.03 % of Topic 2, and less than 5 % of other topics. The student wrote "By setting the system to everything we can make the equation equal to zero. From here, all we need to do is identify everything in the equation, plug in the relating equations, and simplify." This student does not discuss any physics principles to any degree and focuses on their mathematical steps, which is in agreement with our classification of Topic 8 consisting of essays focusing on procedure.
Second, we looked at Spring 2023 in Figure 2(b). We found the topics to be fairly evenly distributed with slight valleys at Topics 4, 5, 6, and 9. Our overall takeaway from these representative words is that there was some focus on explicit physics principles (i.e. Topics 0, 3, 4, and 6), shown from emergent representative words such as "energy", "force", "particle", and "work". An exemplar essay from a student belonging to Topic 3 is Friction force is reliant on the normal force. Because the weight doesn’t change and an additional force is not applied, the normal force does not change.. Though the student is not focusing on energy, they are clearly thinking about other physics involved in solving the problem, such as friction. The other topics there was much more focus on mathematical steps (i.e. Topics 2, 5, 7, and 9), which we see from the representative words such as "used", "equation", and "solve", and surface features (i.e. Topic 1), which we see from the representative words such as "cylinder" and "incline".
Third, we looked at Fall 2023, shown in Figure 2(c). We found the topics to be not as evenly distributed with a large peak emerging at Topic 1. Our overall takeaway from these representative words is that there was some focus on the energy physics principle (i.e. Topics 0, 1, and 4), as words such as "energy", "kinetic", and "extended" are present amongst the top words of these topics. Note that the topic that has the most student membership is Topic 1 which focused primarily on energy. The student essay most representative of Topic 1 is "the total work of the system is equal to translational kinetic evergy plus rotational kinetic energy. This is beacause no other types of energy were gained or lost during this phase. This means we can equal net work to the sum of the energies." This is an exemplar student essay as the student constructs an argument centered around energy principles, such as relating total work to the relevant energy, and conservation of energy. Other topics within this semester that focused on other physics concepts were Topics 2, 3, and 5 where words such as "momentum", "friction", and "force" appear. In addition, the other topics focused on mathematical/procedural steps (i.e. Topic 9), evident from words such as "used" and "find", and surface features (i.e. Topics 6, 7, and 8) concluded from words such as "cylinder", "disk", and "object".
Finally, we looked at Spring 2024 in Figure 2(d). We found the topics to be not as evenly distributed with a large peak emerging at Topic 2. Our overall takeaway from the representative words is that there was some focus on the correct physics principle (i.e. Topics 0 and 2) where words such as "particle" and "energy" emerged. Note that the topic that hand the most student membership is Topic 2 which we tagged as focusing primarily on energy due to the representative words. The student essay that was most representative of Topic 2 was "We used the energy principle to in which the change in energy is equal to the work done by the surroundings, work, which is equal to the net force times d. We are informed by the problem that there will be a change in translational kinetic energy and no change in rotational kinetic energy. Since we are informed of that there is a change in translational kinetic energy and no change in rotational kinetic energy we can set change in Ktrans equal to the work done. After identifying the forces acting on the system and summing them, we get the net force times the displacement on the right side of the equation and the change in translational kinetic energy on the left side. We plug in our representative variables and then solve for the Vcm.". Similar to the Fall 2023 student response, this student focused on energy principles as well, shown by their discussion of the relationship between work and energy. This response also contained elements (though significantly less) of Topic 9 which focused in part on mathematical steps and energy. Other topics that focused on other physics concepts were Topics 3, 5 and 6 illustrated by the representative words, such as "momentum", "velocity", and "force". There were still topics that focused on mathematical steps (i.e. Topic 9), with words like "solve" and "part" emerging, surface features (i.e. Topics 1 and 4), shown by words such as "cylinder" and "ramp" , and a mixture of these surface features, mathematical steps, and physics principles (i.e. Topics 5, 6, 7, and 8).
[…]
@@
5 Discussion
After investigating the top words and topic distribution, we observed that in Fall 2022 and Spring 2023, during which minimal argumentation scaffolding was provided, the topics seemed to be evenly distributed. These results suggest that in general students focused equally on relevant concepts as much as on trivial concepts, such as mathematical steps.
In Fall 2023 and Spring 2024, in which students received more extensive scaffolding, we saw unequal distributions, with peaks in topics that emphasized physics concepts (e.g. energy). This trend indicates that students in the later semesters tended to focus more on physics-oriented ideas. Ideally, students should focus on the underlying physical principles while constructing arguments, as this reflects conceptual reasoning. Conversely, focusing primarily on mathematical steps reflects an emphasis on procedural aspects of problem-solving.
To further investigate the distribution across topics, we conducted a concentration analysis and report on the concentration factor in each semester in Figure 3. Overall, we found an increase in concentration from Fall 2022 to Spring 2024. The Fall 2023 and Spring 2024 semesters, in which students received the most training in argumentation, have higher concentrations of essays than Fall 2022 and Spring 2023 semesters. According to Sadler et al. [sadler2016understanding], when students achieve a more consistent understanding of the underlying concepts, their responses tend to exhibit less variability. Therefore, the observed increase in concentration may reflect more conceptually coherent and consistent reasoning among students in the later semesters.
Taken together, these findings support two key outcomes. First, the increase in concentration indicates that students’ arguments focus more on the topical physics principles rather than procedural details. Second, since these concentration factors were calculated based upon essay membership across topics, this result also demonstrates that topic modeling successfully captured broad shifts in students’ argumentation patterns across semesters.
It is also interesting to note an emergent pattern between fall and spring semesters. The concentration of Spring 2023 is less than that of Fall 2022, and the concentration of Spring 2024 is less than that of Fall 2023. This recurring pattern suggests a potential difference between students enrolled in fall versus spring semesters. Specifically, students in the fall semesters may demonstrate a stronger ability to construct arguments grounded in physics principles.
One possible explanation lies in the known differences of student populations in fall and spring semesters. Students enrolled in the fall tend to enter the course with a stronger physics and calculus background from high school coursework. In contrast, students in the spring generally have less prior physics experience and are more likely to be required to take a prerequisite calculus course.
To investigate if this pattern extends beyond the topic modeling results, we examined scores from the Energy and Momentum Conceptual Survey (EMCS) [singh2003] administered at the beginning of each semester in the laboratory section of the course. We calculated the average scores for each group of students, shown in Figure 4. The same pattern emerged: students in the spring semesters scored lower on average on the EMCS pre-test than those in the fall semesters. This indicates that students in the Spring semesters had lower prior knowledge than those in the Fall semesters, which is consistent with a greater variability in argumentation quality in the spring semester as shown by the lower concentration factor in the spring semester argumentation data.
This convergence between topic modeling results and independent conceptual measures lends further validity to the concentration factor as an indicator of conceptual focus. Moreover, it suggests that the observed differences between semesters are not merely artifacts of the topic modeling process, but may instead reflect genuine differences in students’ prior preparation and conceptual understanding.
Figure 4: EMCS Average Scores at Start of the Semester